if the length of a rectangle is decreased by twice the width the result is 8cm the perimeter of the rectangle is 40cm find the dimensions
Answer by waynest(281)

length - 2w = 8
length - 2w + 2w = 8 + 2w
length(L) = 2w + 8
width = w
p = 2w + 2L
2w + 2(2w + 8) = 40
2w + 4w + 16 = 40
6w + 16 - 16 = 40 -16
6w = 24
6w/6 = 24/6
w = 4 cm
length = 2(4) + 8 = 16 cm
check:
2(4) + 2(2(4) + 8) = 40
8 + 2(8 + 8) = 40
8 + 2(16) = 40
8 + 32 = 40
40 = 40